OpenStudy (kimo_indie):

Solve the system of linear equations below. x + y = 4 2x + 3y = 0

OpenStudy (anonymous):

x+y=4 MULTIPLY BY -2 & ADD. 2x+3y=0 -2X-2Y=-8 ------------------------- Y=-8 ANS. 2X+3*-8=0 2X-24=0 2X=24 X=24/2 X=12 ANS. PROOF: 12-8=4 4=4

OpenStudy (anonymous):

Multiply first equation by 2, then subtract first equation from 2nd equation, you will get an equation in y, then solve for y.